The textures package^{∗}

Sebastian Rahtz Sebastian.Rahtz@tex.ac.uk 1994/10/02

∗This file has version number v1.5i, last revised 1994/10/02.

February 4, 1995

1 Introduction

This file contains the external font information needed to load the standard PostScript fonts, and suitable package files to utilize them.

From this file all .fd files (font definition files) for the PostScript fonts with old encoding (OT1) are created, with the PostScript names for fonts used by Textures.

2 Customization

The usual customization will be to add support for additional combinations in this file, such as a derived bold slanted Times.

3 The docstrip modules

4 The font definition files

As always we begin by identifying the latest version of the files in the log file. The explicit spaces are necessary in an .fd file and the \string guards against situations where ` is active.

1 ⟨∗fd⟩

 2 \wlog{Fontdef\space file\space produced\space from:\space
 3 \string`textures.dtx'\space \fileversion\space(\filedate)\space(SPQR)}

4 ⟨/fd⟩

 5 %
 6 % \subsection{Old \TeX-encoded Adobe AvantGarde}

7 ⟨∗OT1AvantGarde⟩

 8 \typeout{File \space OT1pag.fd\space loading \space Adobe\space AvantGarde}%
 9 \DeclareFontFamily{OT1}{pag}{}%
10 \DeclareFontShape{OT1}{pag}{m}{n}{<->AvantGarde}{}%
11 \DeclareFontShape{OT1}{pag}{m}{it}{<->ssub * pag/m/sl}{}%
12 \DeclareFontShape{OT1}{pag}{m}{sl}{<->AvantGardeI}{}%
13 \DeclareFontShape{OT1}{pag}{b}{n}{<->AvantGardeB}{}%
14 \DeclareFontShape{OT1}{pag}{b}{sl}{<->AvantGardeBI}{}%
15 \DeclareFontShape{OT1}{pag}{b}{it}{<->ssub * pag/b/sl}{}%

16 ⟨/OT1AvantGarde⟩

17 %

18 % \subsection{Old \TeX-encoded Adobe Bookman}

19 %

20 ⟨∗OT1Bookman⟩

21 \typeout{File \space OT1pbk.fd\space loading \space Adobe\space Bookman}%
22 \DeclareFontFamily{OT1}{pbk}{}%
23 \DeclareFontShape{OT1}{pbk}{m}{n}{<->Bookman}{}%
24 \DeclareFontShape{OT1}{pbk}{m}{it}{<->BookmanI}{}%
25 \DeclareFontShape{OT1}{pbk}{b}{n}{<->BookmanB}{}%
26 \DeclareFontShape{OT1}{pbk}{b}{it}{<->BookmanBI}{}%

27 ⟨/OT1Bookman⟩

28 % 29 %

30 % \subsection{Old \TeX-encoded Adobe Courier}

31 %

32 ⟨∗OT1Courier⟩

33 \typeout{File \space OT1pcr.fd\space loading \space Adobe\space Courier}%
34 \DeclareFontFamily{OT1}{pcr}{\hyphenchar \font\m@ne}%
35 \DeclareFontShape{OT1}{pcr}{m}{n}{<->Courier}{}%
36 \DeclareFontShape{OT1}{pcr}{m}{sl}{<->CourierI}{}%
37 \DeclareFontShape{OT1}{pcr}{b}{n}{<->CourierB}{}%
38 \DeclareFontShape{OT1}{pcr}{b}{sl}{<->CourierBI}{}%
39 \DeclareFontShape{OT1}{pcr}{m}{it}{<->ssub * pcr/m/sl}{}%
40 \DeclareFontShape{OT1}{pcr}{b}{it}{<->ssub * pcr/b/sl}{}%

41 ⟨/OT1Courier⟩

42 %

43 % \subsection{Old \TeX-encoded Adobe Helvetica}

44 %

45 ⟨∗OT1Helvetica⟩

46 \typeout{File \space OT1phv.fd\space loading \space Adobe\space Helvetica}%
47 \DeclareFontFamily{OT1}{phv}{}%
48 \DeclareFontShape{OT1}{phv}{m}{n}{<->Helvetica}{}%
49 \DeclareFontShape{OT1}{phv}{m}{it}{<->ssub * phv/m/sl}{}%
50 \DeclareFontShape{OT1}{phv}{m}{sl}{<->HelveticaI}{}%
51 \DeclareFontShape{OT1}{phv}{b}{n}{<->HelveticaB}{}%
52 \DeclareFontShape{OT1}{phv}{b}{it}{<->ssub * phv/b/sl}{}%
53 \DeclareFontShape{OT1}{phv}{b}{sl}{<->HelveticaBI}{}%

54 ⟨/OT1Helvetica⟩

55 %

56 % \subsection{Old \TeX-encoded Adobe New Century Schoolbook}

57 %

58 ⟨∗OT1Schoolbook⟩

59 \typeout{File \space OT1pnc.fd\space loading \space Adobe\space NewCenturySchoolbook}%
60 \DeclareFontFamily{OT1}{pnc}{}%
61 \DeclareFontShape{OT1}{pnc}{m}{n}{<->Schoolbook}{}%
62 \DeclareFontShape{OT1}{pnc}{m}{it}{<->SchoolbookI}{}%
63 \DeclareFontShape{OT1}{pnc}{b}{n}{<->SchoolbookB}{}%
64 \DeclareFontShape{OT1}{pnc}{b}{it}{<->SchoolbookBI}{}%

65 ⟨/OT1Schoolbook⟩

66 %

67 % \subsection{Old \TeX-encoded Adobe Palatino}

69 ⟨∗OT1Palatino⟩

70 \typeout{File \space OT1ppl.fd\space loading \space Adobe\space Palatino}%
71 \DeclareFontFamily{OT1}{ppl}{}%
72 \DeclareFontShape{OT1}{ppl}{m}{n}{<->Palatino}{}%
73 \DeclareFontShape{OT1}{ppl}{m}{it}{<->PalatinoI}{}%
74 \DeclareFontShape{OT1}{ppl}{b}{n}{<->PalatinoB}{}%
75 \DeclareFontShape{OT1}{ppl}{b}{it}{<->PalatinoBI}{}%

76 ⟨/OT1Palatino⟩

77 % \subsection{Old \TeX-encoded Adobe Times Roman}

78 ⟨∗OT1Times⟩

79 \typeout{File \space OT1ptm.fd\space loading \space Adobe\space Times Roman}%
80 \DeclareFontFamily{OT1}{ptm}{}%
81 \DeclareFontShape{OT1}{ptm}{m}{n}{<->Times}{}%
82 \DeclareFontShape{OT1}{ptm}{m}{it}{<->TimesI}{}%
83 \DeclareFontShape{OT1}{ptm}{b}{n}{<->TimesB}{}%
84 \DeclareFontShape{OT1}{ptm}{b}{it}{<->TimesBI}{}%

85 ⟨/OT1Times⟩

86 %

87 % \subsection{Old \TeX-encoded Adobe  Chancery}

88 ⟨∗OT1Chancery⟩

89 \typeout{File \space OT1pzc.fd\space loading \space Adobe\space Chancery}%
90 \DeclareFontFamily{OT1}{pzc}{}%
91 \DeclareFontShape{OT1}{pzc}{m}{it}{<->Chancery}{}%
92 \DeclareFontShape{OT1}{pzc}{m}{n}{<->ssub * pzc/m/it}{}%
93 \DeclareFontShape{OT1}{pzc}{m}{sl}{<->ssub * pzc/m/it}{}%
94 \DeclareFontShape{OT1}{pzc}{b}{n}{<->ssub * pzc/m/it}{}%
95 \DeclareFontShape{OT1}{pzc}{b}{it}{<->ssub * pzc/m/it}{}%
96 \DeclareFontShape{OT1}{pzc}{b}{sc}{<->ssub * pzc/m/n}{}%
97 \DeclareFontShape{OT1}{pzc}{m}{sc}{<->ssub * pzc/m/n}{}%
98 \DeclareFontShape{OT1}{pzc}{b}{sl}{<->ssub * pzc/b/n}{}%

99 ⟨/OT1Chancery⟩

100 %

101 % \subsection{Adobe Symbol font}

102 ⟨∗USymbol⟩

103 \typeout{File \space Upsy.fd\space loading \space Adobe\space Symbol}%
104 \DeclareFontFamily{U}{psy}{}%
105 \DeclareFontShape{U}{psy}{m}{n}{<->Symbol}{}%
106 \DeclareFontShape{U}{psy}{m}{i}{<->ssub * psy/m/n}{}%

107 ⟨/USymbol⟩

108 % \subsection{Adobe Dingbats}

109 ⟨∗UDingbats⟩

110 \typeout{File \space Upzd.fd\space loading \space Adobe\space Dingbats}%
111 \DeclareFontFamily{U}{pzd}{}%
112 \DeclareFontShape{U}{pzd}{m}{n}{<->Dingbats}{}%

113 ⟨/UDingbats⟩